I'm having difficulty determining the following: How would I find the volume and surface area of a cylinder when the radius is 6 and the height is 15?
Thanks!

Answer by checkley71(8403)

THE VOLUME FOR A RIGHT CIRCULAR CYLINDER IS V=pieR~2H or V=3.14*6*6*15 OR
V=1,695.6
THE SURFACE AREA FORMULA IS SA=2pieR~2+2pieRH OR SA=2*3.14*6*6+2*3.14*6*15 or
SA=226.08+282.6 OR SA=508.68